### v1

@@ 4.3 Ablation study
Effectiveness of each component. To demonstrate the effectiveness of each component, we perform detailed ablation experiments in Table 3 and Table 4. In Table 3, the first row shows the results of a frozen pre-trained model without any learnable adapters, while the second row refers to the performance of full fine-tuning. Comparing the results of the third and fourth rows, we can see the effectiveness of PEFT over traditional full fine-tuning in few-shot action recognition. Note that the "Partial Adapting" in the fourth column of Table 3 refers to a common strategy which only adds adapters in the last several ViT layers. We can observe from the table that the improvement of introducing task-specific adaptation into AIM is distinct (1.9%) when inserting adapters into all the layers of the pre-trained model. Additionally, using partial adapting strategy for AIM also brings an improvement of 1%. Even so, the improvement of introducing task-specific adaptation into AIM is notably enhanced under the partial adapting strategy, increasing from 1.9% to 5.9%. In Table 4, the first row presents the baseline result, while the second row shows the performance when applying the three sub-actions, which yields a 1.8% improvement. The third row reports the effect of applying cross attention alignment, leading to a slight decline of 1.7%. Notably, performing an alignment operation without enhancing semantic features is redundant, as the original Task-Adapter already accomplishes semantic alignment during feature extraction. This observation in the fourth row supports our assumption, where the model’s performance increases further from 1.8% to 2.3%. Incorporating order adaptation improves the performance of the Three Sub-actions. It models the sequential relationships between sub-actions in more detail. This change increases performance from 1.8% to 2.5% as shown in the fifth row. The final row summarizes the complete configuration.
Figure 8: Effect of inserting Task-Adapters into the last $L$ ViT layers (e.g., $L$ = 1, 2, 3, 6, 12) on scene-related datasets. The best is highlighted with a red star.
@@
Figure 9: Effect of inserting semantic order adapters into the last $M$ ViT layers (e.g., $M$ = 1, 2, 3, …, 12) on 3 different datasets. The best is highlighted with a black star.
@@
methods visual only semantic only both
@@
Figure 10: We demonstrate accuracy for each time step computed by a single sub-action description to prove its rationality.
@@
The effect of inserting position of adapters. We can see from the Fig. 8 that inserting our Task-Adapter into the top 2 layers is the best setting for all these three datasets. Since these scene-related datasets focus on background understanding and do not need much temporal modeling, a more lightweight setting with inserting adapters into only 2 top layers is more appropriate. In the Fig. 9, we demonstrate the impact of the insertion position of a semantic order adapter. Our analysis indicates that for datasets such as SSv2, which already encode rich action information within their original labels, employing a deep architecture for text sequence modeling is generally unnecessary. We posit that the labels in SSv2, characterized by an abundance of verbs and positional terms, combined with the enhanced contextual information from LLM, yield comprehensive representations. Conversely, for datasets with more straightforward action content, such as Kinetics and HMDB51, a more sophisticated modeling approach is warranted.
The result that using different modalities of information on SSv2-Small 5-way 1-shot. As shown in Table 5, we observe that Task-Adapter++ can significantly improve the classification accuracy of the semantic branch by 3.9%. This improvement is undoubtedly due to the proposed augmentations of semantic information. Our method effectively fills the gap of insufficient information in semantic visual matching, making the two legs of the model equally long.
The accuracy for each frames. In general, LLM provides three descriptions of sub-actions. Except for our claimed method in Section 3, it is easy to consider them as a whole. This actually includes two ideas: averaging three semantic features or concatenating the descriptions together as one prompt. Therefore, Fig. 10 shows the accuracy with different baselines, i.e., without using sub-actions description, averaging features and concatenating three descriptions together. Following Section 3, we separately compute accuracy for each temporal moment using the 3 sub-action descriptions. The results not only indicate that describe sub-actions in this manner is reasonable and helpful, but explain the information loss when just averaging or concatenating them.
$N$-way 1-shot results. Fig. 11 shows that our approach, compared to the original Task-Adapter, not only improves absolute accuracy but also better adapts to $N$-way 1-shot learning ($N$geq 5), i.e., experiences less relative decline. As $N$ increases, classification difficulty also increases, and model performance decline is inevitable. However, because our approach fully considers both semantic and visual features, the model will be less affected in SSv2-Small. And in Kinetics, the results are nearly saturated. It can be seen that the results descend slightly when $N$ increases, and the gap keeps very small.
Figure 12: The choice of different alignment modules on SSv2-Small 5-way 1-shot.
The exploration of different alignment modules. In Fig. 12, we explore different modules for the alignment task. We utilize multi-layer cross attention and 1D CNN (layer=1, stride=1, kernel=3, channel=512) to align visual features. The results indicate that stacking multi-layer cross attention is ineffective and does not improve results, while 1D CNN shows no advantages over cross attention.
@@
Different fine-tuning strategies using different metric measurement methods. We also compare the performance of different fine-tuning strategies using different metric measurement methods to demonstrate the applicability of our Task-Adapter++. Four commonly used methods are taken in this paper, namely TRX [30], Bi-MHM [43], OTAM [4] and ProtoNet (e.g., directly averaging the frame-to-frame cosine similarity over temporal dimension without complex metric measurements). Taking full fine-tuning strategy as the baseline, we observe a consistent improvement across all metric measurement methods, as shown in Fig. 13. It should be noted that ProtoNet outperforms all other methods when uniformly using Task-Adapter++. We attribute this to the fact that with our Task-Adapter++, the model significantly improves the utilization rate of multimodal information and enhances fault tolerance on the basis of Task-Adapter.
Table 6: Effect of inserting position of task adapters and order adapters on SSv2-Small.
@@
Dual Partial Adapting. Firstly, we study the effect of different inserting positions of our Task-Adapter in Table 6 left. The default setting is inserting adapters into all of the 12 ViT layers. Alternatively, we also evaluate the settings which respectively insert adapters to the bottom 2/6 layers and top 2/6 layers of the model (assuming forward propagation is from bottom to top). Inserting adapters only into the top 6 layers has been shown to achieve comparable performance with inserting to all layers in AIM [54]. However, inserting our Task-Adapter only into the top 6 layers outperforms inserting it into all layers by a large margin (5%) for few-shot action recognition. We attribute this to the hypothesize that bottom layers focus on generic features that do not require extensive adaptation, while the top layers focus on task-specific and discriminative features that benefit a lot from our task-specific adaptation module. Additionally, We study the effect of different inserting positions of our semantic order adapter in Table 6 right. The default setting is inserting adapters into all of the 12 ViT layers. We evaluate settings inserting adapters to bottom 2/6 and top 2/6 layers. Notice that inserting adapters only top 2 layers achieves best performance with only 263 K parameters. Considering little difference in results at bottom 2/6 layers, we believe the semantic features have not been well extracted yet, which is exactly the opposite at the top 2/6 layers.
@@
Position of the Task-MSA and O-MSA inner ViT layer. In the visual branch, as illustrated in Section 3, the original AIM comprises two multi-head self-attention (MSA) modules, individually for spatial and temporal modeling, and a task adapter. We study the optimal placement of our proposed task-specific MSA (Task-MSA) relative to the existing two MSAs for improved task modeling. As shown in Table  7 left, inserting the Task-MSA after the temporal (T-) and spatial (S-) MSAs achieves the best performance. It is noteworthy that the Task-MSA consistently outperforms the baseline AIM [54], regardless of its insertion position. These results indicate that the introduction of the Task-MSA indeed enhances few-shot action recognition performance compared to AIM. Furthermore, placing the Task-MSA after the T- and S-MSAs enables maximum extraction of task-specific discriminative features. In the semantic branch, we retain the original Task-Adapter as our baseline without changes to the encoder. We find that relocating the order MSA (O-MSA) back of the original MSA yields higher scores than placing it frontally, as shown in Table 7 right. This is because we model the order relationship of sub-action sequences at the feature level, so positioning the O-MSA later enables better feature-level modeling. Nonetheless, both placements perform better than the baseline Task-Adapter, proving the effectiveness of the Order-MSA.
@@ 5 Conclusion
In this paper, we propose a novel method named Task-Adapter++ to better adapt the pre-trained image models for few-shot action recognition. Instead of using the traditional full fine-tuning strategy, we propose a parameter-efficient dual adaptation method that only fine-tune the newly inserted adapters which alleviates the issues of catastrophic forgetting and overfitting. Specifically, we design task-adapter module to enhance the most discriminative features within the given few-shot learning tasks in visual branch. We otherwise utilize semantic order adapter to model the sequence of sub-actions in semantic branch. We further apply cross attention for cross modal alignment operations, which remarkably improves the performance on five few-shot action recognition benchmarks and achieves the new state-of-the-art results, particularly on the challenging SSv2 dataset.
@@ 6 Acknowledgements
xxx

### v2

@@ 4.3 Ablation study
Effectiveness of each component. To demonstrate the effectiveness of each component, we perform detailed ablation experiments. The ablation study of our Task-Adapter is show in Table 3 and Task-Adapter++ is shown in Table 4. In Table 3, the first row shows the results of a frozen pre-trained model without any learnable adapters, while the second row refers to the performance of full fine-tuning. Comparing the results of the third and fourth rows, we can see the effectiveness of PEFT over traditional full fine-tuning in few-shot action recognition. Note that the “Partial Adapting” in the fourth column of Table 3 refers to a common strategy which only adds adapters in the last several ViT layers. We can observe from the table that the improvement of introducing task-specific adaptation into AIM is distinct (1.9%) when inserting adapters into all the layers of the pre-trained model. Additionally, using partial adapting strategy for AIM also brings an improvement of 1%. Even so, the improvement of introducing task-specific adaptation into AIM is notably enhanced under the partial adapting strategy, increasing from 1.9% to 5.9%.
In Table 4, the first row presents the result of Task-Adapter, while the second row shows the performance when applying the three sub-actions for textual augmentation, which yields a 1.8% improvement. The third row reports the effect of applying cross-modal alignment with fine-grained temporal modeling, leading to a decline of 1.7%. Notably, performing the fine-grained alignment operation without enhancing semantic features is redundant, as the original Task-Adapter already accomplishes semantic alignment during feature extraction. The result in the fourth row supports our assumption, where the model’s performance increases further from 1.8% to 2.3%. Incorporating order adaptation improves the performance of the “Three Sub-actions”. It models the sequential relationships between sub-actions in more detail. This change increases the performance from 1.8% to 2.5% as shown in the fifth row. The last row represents the performance of the complete model, revealing the complementarity and effectiveness of each component.
@@
Figure 8: Effect of inserting Task-Adapter into the last $L$ layers (a) and semantic order adapter (b) into the last $M$ layers (e.g., $L,M$ = 1, 2, 3, …, 12) on 3 different datasets. The best is highlighted with a black star.
The effect of inserting position of adapters. In this part, we explore the result of different inserting positions. We can see from Fig. 8 (a) that inserting our Task-Adapter into the top 2 layers is the best setting for Kinetics and HMDB51. Since these scene-related datasets focus on background understanding, a more lightweight setting with inserting adapters into only the top 2 layers is more appropriate. Inserting the Task-Adapter into the top 6 layers yields the highest performance on SSv2. Since SSv2 requires more intensive temporal modeling than scene-related datasets, a deeper adapter is necessary.
In Fig. 8 (b), we demonstrate the impact of the insertion position of the semantic order adapter. Our analysis indicates that for datasets such as SSv2—which already encode rich action information in their original labels (e.g., “Pouring something into something”)—employing a deep architecture for text sequence modeling is generally unnecessary. Conversely, for datasets with more straightforward action content, such as Kinetics and HMDB51, a more sophisticated modeling approach is warranted.
In conclusion, for scene-related datasets like UCF101, Kinetics and HMDB51, we choose 2-layer Task-Adapter and 8-layer semantic order adapter. For SSv2, we choose 6-layer Task-Adapter and 2-layer semantic order adapter. A detailed analysis of the tunable parameters will be presented in the following subsection.
@@
methods video matching score video-text matching score both
@@
The result of different modalities of information. As shown in Table 5, the integration of multimodal information into the FSAR task significantly enhances the performance of both Task-Adapter and Task-Adapter++. However, we can observe that Task-Adapter achieves an accuracy of only 50.4% when relying solely on the video-text matching score, suggesting that semantic information is not effectively leveraged. While, Task-Adapter++ can significantly improve the classification accuracy of the semantic branch by 3.9%. This improvement can be attributed to the proposed order adaptation of semantic information and fine-grained cross-modal matching. Our method effectively addresses the lack of informative cues in video-text matching, balancing the two branches of the model and enhancing their complementary strengths.
Table 6: The results of different utilization methods for enhanced sub-action description.
methods w/o sub-actions average features concat sub-actions ours
accuracy 60.2 60.7 61.4 63.6
Different methods for using the sub-actions. In general, LLMs provides three descriptions of sub-actions. Except for our claimed method in Section 3, it is easy to consider them as a whole. This actually includes two ideas: averaging the three semantic features or concatenating the descriptions together as one prompt. Therefore, Table 6 shows the accuracy with different strategies, i.e., without using sub-actions description, averaging features and concatenating descriptions together. The results not only demonstrate the effectiveness of decomposing actions into sub-actions, but also reveal the information loss caused by averaging and concatenation, thereby validating the rationale behind stage-wise matching
@@
$N$-way 1-shot results. Fig. 9 shows that our approach, compared to the original Task-Adapter and the baseline, not only improves absolute accuracy but also better adapts to $N$-way 1-shot learning ($N$geq 5), i.e., experiences less relative decline. As $N$ increases, classification difficulty also increases, and model performance decline is inevitable. However, because our approach fully considers both semantic and visual features, its performance is less affected in SSv2-Small. And in Kinetics, the results are nearly saturated. It can be seen that the results descend slightly when $N$ increases, and the gap keeps very small.
Table 7: The results of different alignment modules on 5-way 1-shot task.
Method setting SSv2-Small Kinetics
1-shot 63.6 96.2
1-layer Cross Attnetion 3-shot 69.1 97.3
5-shot 73.2 97.5
1-shot 60.4 94.3
2-layer Cross Attnetion 3-shot 66.7 96.1
5-shot 71.7 97.1
1-shot 60.6 93.3
3-layer Cross Attnetion 3-shot 67.0 95.7
5-shot 71.4 96.8
1-shot 55.4 92.3
Conv1d 3-shot 62.9 94.9
5-shot 68.5 96.0
The exploration of different alignment modules. In Table 7, we explore different temporal modeling modules for the alignment task. We utilize multi-layer cross attention and 1D CNN (layer=1, stride=1, kernel=3, channel=512) to align visual features. The results indicate that stacking multi-layer cross attention is ineffective and does not improve results, while 1D CNN shows significant disadvantage over cross attention.
@@
Different fine-tuning strategies using different metric measurement methods. We also compare the performance of different fine-tuning strategies using different metric measurement methods to demonstrate the applicability of our Task-Adapter++. Four commonly used methods are taken in this paper, namely TRX [15], Bi-MHM [16], OTAM [14] and ProtoNet (e.g., directly averaging the frame-to-frame cosine similarity over temporal dimension without complex metric measurements). Taking full fine-tuning strategy as the baseline, we observe a consistent improvement across all metric measurement methods, as shown in Fig. 10. It should be noted that ProtoNet outperforms all other methods when uniformly using Task-Adapter++. We attribute this to the fact that with our Task-Adapter++, the model significantly improves the utilization of multimodal information and enhances fault tolerance on the basis of Task-Adapter.
Table 8: The tunable parameter analysis of task adapters and order adapters on SSv2-Small.
@@
The tunable parameter analysis. Firstly, we study the accuracy and tunable parameter of different inserting positions of our Task-Adapter in Table 8 left. On SSv2-Small, inserting our Task-Adapter only into the top 6 layers outperforms inserting it into all layers by a large margin (5%) with only 7.2 M parameters. We attribute this to the hypothesize that bottom layers focus on generic features that do not require extensive adaptation, while the top layers focus on task-specific and discriminative features that benefit a lot from task-specific adaptation module. Additionally, We study the accuracy and tunable parameter of different inserting positions of our semantic order adapter in Table 8 right. Note that inserting adapters only into the top 2 layers achieves the best performance with only 263 K parameters. Given the small performance gap at the bottom 2/6 layers, we believe that semantic features are not yet well extracted, which is in contrast to the top 2/6 layers where stronger semantic representations are observed.
In conclusion, for scene-related datasets, Task-Adapter++ utilizes only 2.4 M parameters in the visual branch and 1.1 M parameters in the semantic branch, resulting in a total of 3.5 M parameters. In comparison, for temporal-related datasets, the model employs 7.2 M parameters in the visual branch and 263 K parameters in the semantic branch, amounting to 7.5 M parameters in total. These two strategies both significantly reduce the number of parameters compared to full fine-tuning, which requires 149.6 M parameters.
@@
Position of the Task-MSA and O-MSA inner ViT layer. In the visual branch, as illustrated in Section 3, the original AIM comprises two multi-head self-attention (MSA) modules, individually for spatial and temporal modeling. We study the optimal placement of our proposed task-specific MSA (Task-MSA) relative to the existing two MSAs for task-specific modeling. As shown in Table 9 left, inserting the Task-MSA after the temporal (T-) and spatial (S-) MSAs achieves the best performance. It is noteworthy that the Task-MSA consistently outperforms the baseline AIM [5], regardless of its insertion position. These results indicate that the introduction of the Task-MSA indeed enhances few-shot action recognition performance compared to AIM. Furthermore, placing the Task-MSA after the T- and S-MSAs enables maximum extraction of task-specific discriminative features. In the semantic branch, we retain the original Task-Adapter as our image encoder. We find that placing the order MSA (O-MSA) after the original MSA yields higher scores than placing it frontally, as shown in Table 9 right. This is probably because modeling the order relationships of sub-action sequences benefits more from refined representations, which are typically captured in later stages of the network. Nevertheless, both configurations outperform the baseline Task-Adapter, demonstrating the effectiveness of the proposed O-MSA
Figure 11: Visualization of the attention maps and cross-modal matching scores for the “Pouring something into something” action of different methods. We first present the attention maps generated by AIM and Task-Adapter++ to demonstrate the superiority of our Task-Adapter. Subsequently, we illustrate the cross-modal matching scores at various stages, thereby substantiating the effectiveness of the dynamic sub-action description.
4.4 Visualization results
To further demonstrate the effectiveness of our proposed Task-Adapter++, we visualize the attention maps and cross-modal matching scores in Fig. 11. The input frames are sampled from the “Pouring something into something” action in SSv2-Small. Notably, the attention map generated by AIM is less effective at capturing informative regions, such as the carton of milk, compared to that produced by our Task-Adapter++. Furthermore, our method employs a dynamic sub-action description that segments the complete action into three distinct stages—beginning, process, and end. In comparison with the static prompt, our dynamic prompting strategy yields significant improvements in fine-grained cross-modal matching performance.
@@ 5 Conclusion
In this paper, we propose Task-Adapter++ , a novel framework for adapting pre-trained image models to few-shot action recognition. Instead of traditional full fine-tuning, we introduce a parameter-efficient dual adaptation method that exclusively fine-tunes newly inserted adapters, thereby mitigating catastrophic forgetting and overfitting issues. Specifically, we design a task adapter module to enhance discriminative features in the visual domain for given few-shot learning tasks. Concurrently, we generate sub-action descriptions to capture dynamic label information and employ a semantic order adapter to model the sequential structure of sub-actions in the semantic domain. To better align visual and semantic features for cross-modal matching, we further propose a novel fine-grained cross-modal alignment strategy. Experimental results fully demonstrate the effectiveness and superiority of the the proposed Task-Adapter++, which yields significant performance improvements across five few-shot action recognition benchmarks, particularly on the challenging SSv2 dataset.
@@ 6 Acknowledgements
This work is supported by National Natural Science Foundation of China (No. 62376217, 62301434), Young Elite Scientists Sponsorship Program by CAST (No. 2023QNRC001), and the Joint Research Project for Meteorological Capacity Improvement (Grants 24NLTSZ003).